OpenStudy (anonymous):

f(x) = \sqrt[4]{x^2 - 7 x}.

OpenStudy (anonymous):

f(x) = \sqrt[4]{x^2 - 7 x}. I need to find the domain of this.

OpenStudy (anonymous):

do you mean \[\sqrt[4]{x^2 - 7x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what level of math is this? algebra 1, 2, or calc?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

the x^2-7x is what is in the square root, and the question is asking me to find the domain/

OpenStudy (anonymous):

\[(x^2 - 7x)^.25 = 0\]

OpenStudy (anonymous):

sorry..im having problems with the equation editor...but that is meant as 1/4

OpenStudy (anonymous):

so the domain is 1/4?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

raise both sides of the equation to the 4th power to get \[x^2 -7x = 0\] do you see it now....domain? range?

OpenStudy (anonymous):

I don't remember how to get the domain when it is x^2

OpenStudy (anonymous):

domain is x values

OpenStudy (anonymous):

yes so it would be x^2=7x....so what would I do after that?

OpenStudy (anonymous):

almost...let me write it out...

OpenStudy (anonymous):

\[x(x-7) = 0\]

OpenStudy (anonymous):

you have two values x = 0 and x = 7

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so that means that x can't equal 0 or 7 so how would I write that out using an online application?

OpenStudy (anonymous):

you familiar with set notation?

OpenStudy (anonymous):

no I am not

OpenStudy (anonymous):

\[{x \in R: x \le 0 or x \ge 7}\] means that x is in an element of the real numbers where x is less than or equal to zero OR x is greater than or equal to 7

OpenStudy (anonymous):

is it asking for the range as well?

OpenStudy (anonymous):

no just the domain

OpenStudy (anonymous):

then that's it.

OpenStudy (anonymous):

do you need further assistance?

OpenStudy (anonymous):

what is the notation for greater than or equal, same with smaller than or equal

OpenStudy (anonymous):

\[\ge \] is greater than or equal to....reverse the direct for less than or equal to

OpenStudy (anonymous):

it's not working

OpenStudy (anonymous):